1.) Your town is installing a fountain in the main square. If the water is to rise 28.0 m (91.8 feet) above the fountain, how much pressure must the water have as it moves slowly toward the nozzle that sprays it up into the air? Assume atmospheric pressure equal to 100,000. Pa and g=9.81 m/s2.

2.) Your town is installing a fountain in the main square. If the water is to rise 10.0 m (32.8 feet) above the fountain, how much pressure must the water have as it moves slowly toward the nozzle that sprays it up into the air? Assume atmospheric pressure equal to 100,000. Pa and g=9.81 m/s2.
Answer #1

here,

1)

the height , h = 28 m

atmospheric pressure , P0 = 10^5 Pa

the pressure of water , P = P0 + density of water * g * h

P = 10^5 + 1000 * 9.81 * 28 Pa

P = 3.74 * 10^5 Pa

the pressure must be 3.74 * 10^5 Pa

2)

the height , h = 10 m

atmospheric pressure , P0 = 10^5 Pa

the pressure of water , P = P0 + density of water * g * h

P = 10^5 + 1000 * 9.81 * 10 Pa

P = 1.98 * 10^5 Pa

the pressure must be 1.98 * 10^5 Pa